An opera singer in a convertible sings a note at 850 Hzwhile cruising down the highway at 94 km/h . The speed of sound in the air is 343 m/s.

What is the frequency heard by a person standing beside the road in front of the car?

Express your answer to two significant figures and include the appropriate units.

f+ =

========================

Part B

What is the frequency heard by a person standing beside the road behind the car?

Express your answer to two significant figures and include the appropriate units.

f- =
Answer #1

A )

f = actual frequency = 850 Hz

f' = observed frequency

V= speed of sound = 343 m/s

Vs = speed of source of sound = 94 km/h = 26.11 m/s

VL = speed of listener in front = 0 m/s

since the source is moving towards the listener at rest , we use the formula for observed frequency as

f' = (V- VL ) f / (V - Vs)

f' = (343 - 0) (850) / (343 - 26.11)

f' = 920.04 Hz

B)

when the source (car) moves away from the listener, observed frequency is given as

f' = (V- VL ) f / (V + Vs)

f' = (343 - 0) (850) / (343 + 26.11)

f' = 789.87 Hz